# SwiftTry: Fast and Consistent Video Virtual Try-On with Diffusion Models

### v1

@@ Conclusion
In conclusion, we have proposed a novel technique, ShiftCaching, which ensures temporal smoothness across video clips while effectively reducing redundant computations during video inference. This advancement enhances the efficiency and quality of video virtual try-on, making it more practical for real-world applications. Additionally, we have introduced a new dataset, TikTokDress, designed specifically for video virtual try-on. This dataset stands out for its diverse range of backgrounds, complex movements, and high-resolution videos, addressing the limitations of existing datasets and providing a valuable resource for future research in this area.
@@
In this supplementary material, we present additional experimental results and details that could not be included in the main paper due to space constraints. First, we offer a comprehensive overview of the training process for our proposed method. Next, we delve into the specifics of various masking strategies employed in the attention module. We also provide detailed information of creating our TikTokDress dataset. Finally, we showcase additional qualitative results of our method on the VVT and TikTokDress datasets.
@@ Appendix A Training Detail
Optimization. We use Adam (Kingma and Ba 2014) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, a fixed learning rate of $51e-$ 5 for both image-pretraining stage and stage 2. We also utilize Min-SNR Weighting Strategy (Hang et al. 2023) with $\gamma=5.0$. We conduct our training in mixed-precision fp16.
Hardware. All the codes are implemented in Pytorch utilizing Diffusers (von Platen et al. 2022) framework and we only conduct training on a single A100 40GB in both two stages. The training takes around 3 days for stage 1 on VITON-HD (Choi et al. 2021) and around 30 hours for stage 2 on both VVT and TikTokDress. Additionally, the FPS benchmarks reported in the paper are also obtained using the A100 GPU.
Training. In the first stage, we optimize the Main Unet decoder, Garment Unet, and Pose Encoder on VITON-HD (Choi et al. 2021) while other modules are kept fixed. The training resolution at this stage remains at $1024\times 768$, consistent with the original resolution. Then, we sample a clip of 16 frames with resolution $512\times 384$ for each video as input for stage 2. We optimize motion modules only in this stage.
@@ Appendix B Different Masking Attention Mechanisms
Fig. 9 illustrates the different attention masking strategies in our ShiftCaching method. In the Full Attention configuration ($a$), the attention mask is filled with $0.0$, allowing each frame to attend to all others. However, this can lead to less accurate features (i.e., $t+2$, shown in white) influencing the more accurate ones (i.e., $t+1$, shown in red). In the Half Attention ($b$) setup, the query frame always attends to the accurate features ($t+1$), enabling the transfer of quality features to less accurate ones. The Quarter Attention($c$) strategy allows frames with less accurate features to attend to all other frames while ensuring that the accurate features do not attend to less accurate ones, thereby preserving their quality. Finally, Causal Attention ($d$) ensures that the computation for each frame should only depend on the succeeding frames. We perform experiments and find out that Half Attention($b$) yields the best performance.
Figure 9: The attention mask used in different masked temporal attention mechanisms in ShiftCaching. White cells in the matrix represent a value of 0.0, while gray cells indicate a value of $\infty$- ∞.
@@ Appendix C TikTokDress Dataset
Fig. 12 showcases examples from our dataset, which is diverse in background (ranging from indoor to outdoor settings), camera positions (from top to bottom angles), person distances (from far to close to the camera), skin tones (including black, yellow, and white), and garment types (such as t-shirts, sweaters, crop tops, long-sleeve t-shirts, and dresses). Based on this diversity, we believe that our dataset better reflects real-life scenarios for virtual try-ons.
Figure 10: Impact of mask quality on try-on results. The first row shows target clothing images. The second row demonstrates the effects of an inaccurate mask, leading to flawed try-on results in the third row. The fourth row highlights how a precise mask significantly improves quality and accuracy. Accurate masks are crucial for high-quality try-ons.
@@
Agnostic masks are crucial for achieving accurate try-on results. For example, in Fig. 10, an incorrect agnostic mask (highlighted by the red circle) causes the model to hallucinate a garment to fit the mask, resulting in a try-on outcome where the garment appears as a tank top (highlighted by the blue rectangle in the first row). With a correct garment mask, a good try-on result is achieved (highlighted by the green rectangle in the first row). In the second row, a garment that is not fully masked leads to the model hallucinating additional red stripes in the middle of the red garment (highlighted by the blue rectangle in the second row). Again, a correct garment mask produces a good try-on result (highlighted by the green rectangle in the second row).
@@
To obtain accurate agnostic masks, we use SAM2 (Ravi et al. 2024) for both garment and arm segmentation. Given a video with multiple frames showing diverse poses, we select the frame that displays as many visible body parts as possible and is clear, similar to Fig. 11. We apply the algorithm described in Algorithm 1 for this frame selection.
The agnostic masks are a combination of the garment mask and the arm mask. For garment masks, we choose positive points in the prompting process, specifically the body center, right shoulder, and left shoulder. The body center is defined as the midpoint between the neck, right hip, and left hip (see points 1, 8, and 11 in Fig. 11). For the negative points, we select noise, the left eye, and the right eye. For arm masks, the positive points we choose are the left elbow, right elbow, left wrist, and right wrist. For the negative points, we choose the right knee, nose, and left knee. After obtaining the garment and arm masks, we combine them and dilate the result three to seven times to ensure it fully conceals the original garment’s shape.
Garment Image. To extract garment images from videos, we use the Google Lens engine to find the best match at the highest possible resolution. In cases where Google Lens does not return a satisfactory result, we manually search on Google using keywords related to the type of clothing, color, and brand, selecting the best match. If we are unable to find a suitable garment image, we exclude that video from our dataset. After obtaining the garment image, in some cases, we horizontally flip the image to align it with the video.
@@ Appendix D More Visualization Results
In the video try-on task, we compare our method on the VVT dataset (Dong et al. 2019b) against ViViD (Fang et al. 2024) in Fig. 13. As demonstrated in video $b.mp4$(see in Supplemental Videos), our method consistently preserves the garment texture throughout the entire video better than ViViD in maintaining texture stability over time. In video $d.mp4$ (see in Supplemental Videos), the “GAP” letters on the garment in ViViD exhibit noticeable flickering, a result of its temporal averaging technique. In contrast, our method, which employs the ShiftCaching technique, significantly reduces this flickering, ensuring smoother and more stable visual results (Notes: we recommend to see in the supplementary videos).
Figure 13: Qualitative comparison between our method and ViViD. Our model achieved robust results through long video, thereby forming a coherent video sequence.
Figure 14: More examples of our video try-on results on our TikTokDress dataset. The first row illustrates how well our method’s performs in various backgrounds. The second row highlights our method’s ability to preserve texture consistency over time.
@@
Fig. 14 showcases additional results of our video virtual try-on method on the TikTokDress dataset. By utilizing SAM 2 (Ravi et al. 2024) and manually refining masks on our dataset, combining with the SAM2 frame selection algorithm Algorithm 1, our try-on results demonstrate superior performance compared to those trained on the VVT dataset. Our method ensures consistent texture quality for garments and, when combined with ShiftCaching, effectively produces continuous and smooth long videos (Notes: we recommend to see more detail in the supplementary videos).

### v2

@@ Conclusion
In conclusion, we have proposed a novel technique, ShiftCaching, which ensures temporal smoothness across video chunks while effectively reducing redundant computations during video inference. This advancement enhances the efficiency and quality of video virtual try-on, making it more practical for real-world applications. Additionally, we have introduced a new dataset, TikTokDress, designed specifically for video virtual try-on. This dataset stands out for its diverse range of backgrounds, complex movements, and high-resolution videos, addressing the limitations of existing datasets and providing a valuable resource for future research in this area.
@@
In this supplementary material, we present additional experimental results and details that could not be included in the main paper due to space constraints. First, we provide a comprehensive overview of the training process for our proposed method. Next, we elaborate on the various masking strategies employed in the attention module. We also include detailed information about the creation of our TikTokDress dataset. Finally, we showcase additional qualitative results of our method on both the VVT and TikTokDress datasets.
@@ Appendix A Training Detail
Optimization. We use the Adam (Kingma and Ba 2014) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and a fixed learning rate of $51e-$ 5 for both the image pretraining stage and stage 2. We also adopt the Min-SNR Weighting Strategy (Hang et al. 2023) with $\gamma=5.0$. All training is conducted in mixed-precision FP16.
Hardware. The code is implemented in PyTorch using the Diffusers (von Platen et al. 2022) framework, and training is performed on a single A100 GPU with 40GB memory for both stages. The training for stage 1 on VITON-HD (Choi et al. 2021) takes approximately 3 days, while stage 2 on both the VVT and TikTokDress datasets takes around 30 hours. The FPS benchmarks reported in the paper were also measured using the A100 GPU.
Training. In the first stage, we optimize the Main UNet decoder, Garment UNet, and Pose Encoder on VITON-HD (Choi et al. 2021), keeping all other modules fixed. The training resolution for this stage is maintained at $1024\times 768$, consistent with the original dataset resolution. In the second stage, we sample a clip of 16 frames with a resolution of $512\times 384$ from each video as input. At this stage, we optimize only the motion-related modules.
@@ Appendix B Different Masking Attention Mechanisms
Fig. 9 illustrates the different attention masking strategies in our ShiftCaching method. In the Full Attention configuration ($a$), the attention mask is filled with $0.0$, allowing each frame to attend to all others. However, this setup can result in less accurate features (e.g., $t+2$, shown in white) influencing more accurate ones (e.g., $t+1$, shown in red). In the Half Attention setup ($b$), the query frame always attends to the accurate features ($t+1$), enabling the transfer of high-quality features to less accurate ones. The Quarter Attention setup ($c$) allows frames with less accurate features to attend to all others while ensuring that accurate features do not attend to less accurate ones, thereby preserving their quality. Finally, the Causal Attention setup ($d$) ensures that each frame’s computation depends only on succeeding frames (less noise frames). Our experiments reveal that Half Attention ($b$) yields the best performance.
Figure 9: The attention mask used in different masked temporal attention mechanisms in ShiftCaching. White cells in the matrix represent a value of 00, while gray cells indicate a value of $\infty$- ∞.
@@ Appendix C TikTokDress Dataset
Fig. 12 showcases examples from our dataset, highlighting its diversity in backgrounds (ranging from indoor to outdoor settings), camera positions (including top and bottom angles), person distances (ranging from far to close to the camera), skin tones (spanning Black, Asian, and White), and garment types (such as T-shirts, sweaters, crop tops, long-sleeve T-shirts, and dresses). This diversity makes our dataset more representative of real-life scenarios for virtual try-on applications.
Figure 10: Impact of Mask Quality on Try-On Results. The first row displays the target clothing images. The second row illustrates the effects of an inaccurate mask, resulting in flawed try-on outputs shown in the third row. The fourth row demonstrates how a precise mask significantly enhances the quality and accuracy of the results. Accurate masks are essential for achieving high-quality try-on outcomes.
@@
Agnostic masks are essential for achieving accurate try-on results. For instance, in Fig. 10, an incorrect agnostic mask (highlighted by the red circle) prompts the model to hallucinate a garment to fit the mask, resulting in a try-on outcome where the garment appears as a tank top (highlighted by the blue rectangle in the first row). Conversely, using a correct garment mask produces a satisfactory try-on result (highlighted by the green rectangle in the first row). Similarly, in the second row, a garment that is not fully masked causes the model to hallucinate additional red stripes in the middle of the red garment (highlighted by the blue rectangle in the second row). Once again, a correct garment mask leads to an accurate try-on result (highlighted by the green rectangle in the second row).
@@
To obtain accurate agnostic masks, we use SAM2 (Ravi et al. 2024) for both garment and arm segmentation. From a video with multiple frames showing diverse poses, we select the frame that displays the most visible body parts and has the highest clarity, as illustrated in Fig. 11. We apply the algorithm described in Algorithm 1 to facilitate this frame selection.
The agnostic masks are created by combining the garment mask and the arm mask. For garment masks, we select positive points during the prompting process, specifically the body center, right shoulder, and left shoulder. The body center is defined as the midpoint between the neck, right hip, and left hip (refer to points 1, 8, and 11 in Fig. 11). Negative points include noise, the left eye, and the right eye. For arm masks, the positive points selected are the left elbow, right elbow, left wrist, and right wrist, while negative points include the right knee, nose, and left knee. After generating the garment and arm masks, we combine them and apply dilation three to seven times to ensure the mask fully conceals the original garment’s shape.
Garment Image. To extract garment images from videos, we use the Google Lens engine to find the best match at the highest possible resolution. If Google Lens does not provide a satisfactory result, we manually search on Google using keywords related to the garment type, color, and brand, selecting the most suitable match. If no appropriate garment image can be found, we exclude the video from our dataset. After obtaining the garment image, we occasionally horizontally flip it to align with the video.
@@ Appendix D More Visualization Results
In the video try-on task, we compare our method on the VVT dataset (Dong et al. 2019b) with ViViD (Fang et al. 2024), as shown in Fig. 13. In video $b.mp4$ (see Supplemental Videos), our method consistently preserves garment textures throughout the video, outperforming ViViD in maintaining texture stability over time. Similarly, in video $d.mp4$ (see Supplemental Videos), the”GAP” letters on the garment produced by ViViD exhibit noticeable flickering due to its temporal averaging technique. In contrast, our method, which incorporates the ShiftCaching technique, significantly reduces flickering, ensuring smoother and more stable visual results. We recommend viewing the supplementary videos for a detailed comparison.
Figure 13: Qualitative comparison between our method and ViViD. Our model achieves robust results over long videos, producing a coherent video sequence.
Figure 14: Additional examples of our video try-on results on the TikTokDress dataset. The first row illustrates how effectively our method performs across various backgrounds, while the second row highlights its ability to preserve texture consistency over time.
@@
Fig. 14 showcases additional results of our video virtual try-on method on the TikTokDress dataset. By leveraging SAM 2 (Ravi et al. 2024) and manually refining masks in our dataset, combined with the SAM 2 frame selection algorithm (Algorithm 1), our try-on results demonstrate superior performance compared to those trained on the VVT dataset. Our method ensures consistent garment texture quality and, when paired with ShiftCaching, effectively generates continuous and smooth long videos. More details can be found in the supplementary videos).